OpenStudy (anonymous):

What is the acceleration of a cabinet of mass 45 kilograms if Jake and Ted push it by applying horizontal force of 25 newtons and 18 newtons respectively in the same direction? (Assume a flat, frictionless surface.)

OpenStudy (shane_b):

Remember what I told you on the last question: The net force will be the sum of the forces applied. In this case, the forces are in the same direction so they will have the same sign. Therefore, the net force will be:\[25N+18N=43N\]From there, you can solve for acceleration like we did on the last problem.

OpenStudy (anonymous):

we had to divide but i tried to divide the 45 with 43 it didnt make sence or do i switch the two?

OpenStudy (shane_b):

\[F=ma\]Rearrage to get:\[\frac{F}{m}=a\]Put in the values you know:\[\frac{43N}{45kg}=? \space m/s^2\]

OpenStudy (anonymous):

ohhh right okay i see

OpenStudy (shane_b):

You should get:\[a=1 \frac{2}{43}m/s^2\]

OpenStudy (shane_b):

Just about 1 m/s^2 if you round it

OpenStudy (anonymous):

0.9555555 im using my phone for a calculator but thats what i got

OpenStudy (shane_b):

Looks right to me :) If you're supposed to use significant figures the answer would be 0.96 m/s^2

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

yay i passed

OpenStudy (shane_b):

Good, keep practicing :)

OpenStudy (anonymous):

thank you i will